40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 18 02:42:28 2014, Jenne, Update, LSC, AO path partly engaged Zoom_TRXTRY4_EngagingAO_1095059162.png0020.pdf
    Reply  Thu Sep 18 17:44:55 2014, Jenne, Update, LSC, Old AO cable pulled 
       Reply  Tue Sep 30 17:26:18 2014, ericq, Update, LSC, New AO cable in place 
    Reply  Fri Sep 19 04:05:05 2014, ericq, Update, LSC, AO path partly engaged TFSR785_19-09-2014_020555.pdfTFSR785_19-09-2014_033920.pdf
       Reply  Fri Sep 19 13:12:07 2014, Jenne, Update, LSC, AO path glitches 
Message ID: 10554     Entry time: Tue Sep 30 17:26:18 2014     In reply to: 10519
Author: ericq 
Type: Update 
Category: LSC 
Subject: New AO cable in place 

I've installed a new 2pin lemo cable going from the CM servo out to in2 of the MC servo board, and removed the temporary BNC. I used some electrical tape to give the cable some thickness where the lemo head screws on to try to strain relieve the solder joints; hopefully this cable is more robust than the last. 

I put an excitation into the CM board, and saw it come out of MC_F, so I think we're set. 

ELOG V3.1.3-